# Deep Learning based Key Information Extraction from Business Documents: Systematic Literature Review

### v1

@@ Abstract.
Extracting key information from documents represents a large portion of business workloads and therefore offers a high potential for efficiency improvements and process automation. With recent advances in deep learning, a plethora of deep learning-based approaches for Key Information Extraction have been proposed under the umbrella term Document Understanding that enable the processing of complex business documents. The goal of this systematic literature review is an in-depth analysis of existing approaches in this domain and the identification of opportunities for further research. To this end, 96 approaches published between 2017 and 2023 are analyzed in this study.
@@
copyright: rightsretainedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: CSURccs: Applied computing Document analysisccs: Applied computing Business process management systemsccs: Computing methodologies Information extractionccs: Computing methodologies Neural networks
@@ 1. Introduction
Complex business documents pose a significant challenge to KIE systems because they cannot be understood and processed as linear text sequences, as has been the case in most traditional KIE applications (Borchmann et al., 2021). These documents typically contain implicit and explicit cues such as different font formatting to distinguish headers from other parts of the document or complex positional dependencies between certain text segments. In this regard, related research investigates different methods to integrate such cues into the model architectures in order to achieve better extraction results. Business documents also have other special characteristics resulting from their connection to business processes. For example, different documents that are being processed as part of the same process run usually contain reoccurring information. These aspects and how they can improve corresponding KIE systems are worth investigating. In general, a process-oriented understanding is critical in order to adequately address the challenge of KIE in real-world contexts.
@@
    A SLR based on 96 approaches to provide a concise overview of DL-based KIE methods for business documents.
@@
The remainder of this manuscript is structured as follows: Chapter 2 provides background information on key concepts and nomenclature that are relevant to this SLR. In chapter 3, we discuss related work in terms of existing surveys and illustrate, how this study differs. The methodology used for this SLR, and more specifically how relevant literature was identified, is illustrated in chapter 4. Chapter 5 provides an overview of the results of the in-depth analysis. Chapter 6 serves as a discussion of the key findings and is dedicated to specific aspects of the identified literature. Based on the previous two chapters, we propose a research agenda and starting points for follow-up research in chapter 7. Chapter 8 concludes the manuscript with a summary.
@@ 2.1. A business process perspective
In addition to implicit knowledge about predefined entities, the consideration of associated business processes can also be useful in the sense that the information flow between individual process steps could be taken into account. For example, one could consider how the data extracted from a business document is further processed in subsequent process steps. From this, conclusions could be drawn as to how corresponding business documents should be processed. One could also analyze whether, in addition to the documents themselves, there is an additional data flow between the document processing steps that could facilitate the understanding of the documents.
@@ 2.2. Document Understanding
Document Understanding (DU), based on concepts from Natural Language Processing (NLP) and Computer Vision (CV), is an umbrella term covering a wide range of document processing tasks, including KIE, table understanding, document layout analysis (DLA), document classification and visual question answering (VQA) (Borchmann et al., 2021). More recently, novel tasks have been proposed such as ”Key Information Localization and Extraction”, which is an extension of KIE that emphasizes the need to localize key information, e.g., by identifying bounding boxes [^1] A bounding box indicates the coordinates of an element in the document image by its top left corner as well as its width and height. (Skalický et al., 2022). For a brief introduction to DU tasks besides KIE that are not covered in this review, we refer to (Cui et al., 2021). Other terms are also used for this research area, such as Document Analysis, Document AI or Document Intelligence (Motahari et al., 2021; Cui et al., 2021).
@@ 2.4. Input documents
The term ”business document” appears less frequently in related literature and there is also no universal definition of such documents (Cristani et al., 2018). In general, business documents contain process-relevant details related to the internal and external operations of an organization, as these documents represent a central means of communication (Cui et al., 2021; Skalický et al., 2022). Business documents, like VRDs, pose a significant challenge to automated DU systems for a variety of reasons, as discussed by (Motahari et al., 2021). For example, corresponding documents exist in a variety of different formats and are often only available in scanned form due to their paper-based distribution. The layout and overall content of business documents can vary from highly structured to highly unstructured. Furthermore, business documents can have relationships to other documents and/or may consist of multiple documents in a hierarchical fashion. To understand business documents, it is therefore necessary to observe such interrelationships and to understand the temporal relationships in processes.
@@ 3. Existing reviews on Key Information Extraction
In the following, we discuss the most closely related work. (Subramani et al., 2020) provide a brief KIE section that is limited to discussions on a few relevant aspects. The survey also does not include an in-depth analysis of the approaches. In (Baviskar et al., 2021), the authors provide a comprehensive overview of document processing with many different facets. However, it offers limited discussion about KIE. Much refers specifically to NER, where the overall workflow as well as different methods regarding pre-processing and feature extraction are being discussed. The authors discuss some model architectures. Since the chosen search period includes work between 2010 and 2020, a large number of analyzed papers also do not necessarily use DL methods and/or consider VRDs. (Cui et al., 2021) cover a wide range of DL-based DU topics. KIE, on the other hand, is discussed relatively briefly. The survey discusses aspects related to transformer-based document processing methods from a general point of view. However, details about the employed model architectures and other KIE paradigms are not covered in detail. The survey of (Antonio et al., 2022) focuses on preliminary DU tasks such as text detection and text recognition. KIE is also covered, although only a small number of approaches are included. The authors present some individual approaches in detail, however there is a lack of comprehensive overview. In (Martínez-Rojas et al., 2023), document processing is positioned in the context of Robotic Process Automation. Although an extensive literature survey is conducted, there is no in-depth technical discussion of KIE approaches for VRDs. An overview of key methods for KIE is provided by (Li et al., 2023a), albeit at a relatively high level. The authors however provide an extensive outlook on future work. Besides KIE, (Sassioui et al., 2023) also cover VQA and document classification. The authors propose a taxonomy for DU along different dimensions and focus on KIE benchmark datasets. Some challenges as well as future work are also discussed. (Abdallah et al., 2024) present a review of transformer-based methods for DU tasks. They highlight key paradigms of corresponding models and showcase a few approaches in detail. A major focus of the review is the description of benchmark datasets and related performance comparisons. In general, the survey chooses to discuss a few transformer-based approaches in detail, but does not present a broad overview of DL-based approaches with respect to KIE.
@@
Overall, existing surveys examine KIE from a very broad perspective, cover only a few approaches in detail, or alternatively include a higher number of published work, but at the expense of a detailed discussion of the underlying methods. This work on the other hand provides an in-depth analysis, both quantitatively and qualitatively. One key distinguishing factor is also the adoption of a business process perspective. As discussed in section 2.1, the consideration of business processes as well as general domain knowledge is of high importance for KIE systems. To this end, only the review by (Martínez-Rojas et al., 2023) also adopts a practice-oriented perspective during the analysis. Table 1 summarizes the differentiation of the previously mentioned related work along certain distinguishing factors.
@@ 4.1. Research questions
The first five research questions cover an in-depth analysis of the identified approaches, with the aim of examining the proposed methods and how they differ from each other. This includes key aspects such as input modalities, model architectures and data bases. Based on this, research questions 6 to 8 adopt a more aggregated view and aim to identify the state of the art and derive recommendations for follow-up research.
@@ 4.2. Search procedure
The following six databases were used to obtain relevant literature: ACM Digital Library (ACM), ACL Anthology (ACL), AIS eLibrary (AIS), IEEE Xplore (IEEE), ScienceDirect (SD) and SpringerLink (SL). The search strings were carefully designed to include the commonly used terms for the research area as well as relevant keywords for the target domain (VRDs) and the application of DL architectures. If supported by the search engine, we also included terms that should not appear to filter out papers that are outside the scope of this work. The search period was limited to the range 2017 to 2023. Considering 2017 as the lower limit is a significant help to avoid false positive results, as KIE related literature typically did not use DL methods before 2017. The final search strings for each database, including the number of results at the time the query was run, can be found in table ST1 of the electronic supplementary material.
@@
  • EX3: The work does not apply the approach to VRDs, but to other domains or only considers text-only inputs.
@@
The overall search process is visualized in figure 3 according to the PRISMA defined by (Page et al., 2021). A total of 96 relevant approaches were identified for this SLR. A list of the identified approaches as well as their numbering, which is also used as a reference in this manuscript, can be found in table LABEL:mapping of the electronic supplementary material.
Figure 3. PRISMA flow diagram
[…]
Literature search procedure according to PRISMA flow diagram
@@ 5.1. Overview
First, we provide a high-level overview of the analyzed work. To this end, table LABEL:overviewtab of the electronic supplementary material shows the results of the in-depth analysis along different properties, each of which considers different aspects, also with practical applications in mind. Figure 4(a) shows the overall distribution of the KIE paradigms. Almost half of the approaches belong to the group of sequence-based methods (n=46). Two further common categories are graph-based (n=24) and a combination of graph-based and sequence-based approaches (n=10). The grid-based representation is not as widely used as the graph-based representation, as only 9 approaches are purely based on this paradigm. Also, only 6 of the analyzed papers cannot be assigned to any of the groups, which underlines the dominance of the three paradigms for KIE systems. Interestingly, only the work by (Wang et al., 2021b) combines grid-based and sequence-based concepts into one KIE approach. Based on the distribution of the methods over time, as visualized in figure 4(b), one can see an increased interest in KIE research, especially since 2020. It is also noticeable that the dominance of sequence-based approaches first started in 2021, while graph-based methods were the most common group of approaches before that. The increased popularity of sequence-based methods in this period can probably be attributed to the influential work by (Xu et al., 2020), which was among the first to show state-of-the-art results using a transformer architecture and extensive pre-training procedures, which in turn led to a lot of follow-up research in this direction. One can also see a strong increase of graph-based methods in 2023 compared to the previous years as well as compared to the remaining categories. This category has therefore remained relatively popular over time. Overall, 2021 and 2023 saw the highest number of published papers. In 2022, however, there was a decrease in the number of corresponding KIE approaches (18 compared to 27 in the previous year). One reason for this observation may be the aforementioned increased interest in sequence-based methods, which typically require extensive pre-training. Therefore, many scholars spent 2022 on developing corresponding methods and on academic write-up, leading to a publication some time later in 2023.
@@
The analysis of the integrated input modalities shows that the vast majority of approaches integrate at least textual (n=85) and layout-oriented (n=76) features. This is not surprising, since these elements contain the most relevant information for document processing. Layout modalities are usually derived from bounding boxes, i.e., the coordinates of words in document images. 56 of the analyzed approaches also integrate visual features obtained from the document images. As discussed in section 2.4, the integration of such visual cues into KIE systems is crucial for a proper understanding of complex VRDs. In this regard, one could even argue that, given this relevance, there are rather few approaches that integrate image modalities. Nevertheless, the results show that visual cues have became more and more important and have been increasingly integrated into the models over time, especially since 2021. In this regard, (Oral and Eryiğit, 2022) extensively investigate different methods to fuse visual modalities into KIE systems. Integrating multiple input modalities can play an important role for DU systems, as they incorporate cues from different aspects and therefore complement each other. For example, the image modality can provide relevant insights for documents with otherwise limited amounts of text and bounding boxes, while the layout modality is crucial for documents with very complex layouts and distinct structures (Yang et al., 2024). 36 of the approaches use hand-crafted input modalities. In most of these cases, they are integrated through custom input features such as boolean flags, whether a word in the document represents specific entities such as dates or monetary values (Palm et al., 2017; Gal et al., 2019; Lohani et al., 2019; Gal et al., 2020; Belhadj et al., 2021; Hamdi et al., 2021; Holeček, 2021; De Trogoff et al., 2022). Another feature type is information about the position of a particular word with respect to the overall reading order (Lee et al., 2021; Shi et al., 2023). In some papers, hand-crafted inputs are based on syntactic features such as the numbers of characters of a word (Sage et al., 2020; Hamri et al., 2023), details regarding the fonts (Wei et al., 2020; Hwang et al., 2021a) or encoded character representations (Krieger et al., 2021). (Zhang et al., 2023a) propose a feedback mechanism by fusing a label embedding based on decoder outputs back into the encoder for predicting adjacent fields. Therefore, the predicted label of an element also affects the prediction for adjacent elements. The use of hand-crafted features is also often associated with grid-based approaches. Such methods usually define a custom encoding function that assigns a specific value to each element of the grid. For example, (Palm et al., 2017; Dang and Thanh, 2020) assign a constant value to each pixel in a document image, such as an integer index of a character at each position – or the value 0 for blank parts of the document image. None of the approaches integrate meta data resulting from related business workflows as discussed in section 2.1, therefore lacking a practice-oriented perspective in this regard. Overall, hand-crafted features provide additional meta-level information beyond data that is explicitly contained in documents and can thus improve KIE systems. It must be said, however, that hand-crafted input features can potentially lead to a considerable amount of additional work in terms of labor-intensive data annotations (Saifullah et al., 2023a), which could also explain the less frequent use of this feature type. One could also argue that the inclusion of such hand-crafted features can potentially result in overfitting to the domain of the training documents – leading to poorer generalization capabilities across different domains. However, there is no work that analyzes this trade-off between integrating hand-crafted input features and the generalization capabilities across different domains.
@@
In terms of the underlying data basis, the median number of employed documents is 2,702. This is a relatively low amount, especially since the data basis usually has to be split into training and evaluation partitions. As a result, many of the approaches are often only trained on a small document corpus and thus probably with little variety in terms of layouts. Since DL-based DU models usually require large amounts of data for training (Saifullah et al., 2023a), it is debatable whether the proposed approaches have reached their full potential. Note that these observations do not include the data basis that is being used for pre-training purposes, which will be discussed in the context of sequence-based approaches in section 5.2.1. Deviating from the average, (Palm et al., 2019; Klaiman and Lehne, 2021; Sarkhel and Nandi, 2021) employ more than 1 million documents for implementing their KIE system. On the other hand, in some cases, only 199 documents are used for developing the models, while still achieving promising extraction results (Li et al., 2021a, b; Du et al., 2022; Gemelli et al., 2023; Yu et al., 2023). The vast majority of used document types are receipts, forms and invoices. These three categories alone are used by around 70 % of the approaches. The main reason for this is the fact that the most frequently used benchmark datasets, namely CORD, FUNSD and SROIE [^3] These datasets will be discussed in more detail in section 5.3.1., are based on these document types. Invoices, on the other hand, are most often considered when authors do not use public benchmarks, but rather private in-house datasets.
@@
Regarding the aspect of reproducibility, implemented code is available for 38 approaches. In most cases, authors share their code directly, however there are exceptions where implementations have been made available by external sources through re-implementation (e.g., the approach by (Powalski et al., 2021)). Model weights are shared even less frequently, which is however especially helpful for pre-trained sequence-based approaches. When either code or model weights are shared, they are mostly provided with a license that allows their application in commercial settings. This is beneficial for organizations that want to integrate corresponding KIE systems into their own business processes, which in turn helps to disseminate the research efforts.
@@
10 approaches are independent of external OCR engines (Guo et al., 2019; Zhang et al., 2020; Wang et al., 2021a; Gu et al., 2022; Kim et al., 2022; Davis et al., 2023; Dhouib et al., 2023; Kuang et al., 2023; Yang et al., 2023a; Yu et al., 2023) and are therefore either responsible both for text reading and information extraction (end-to-end) or alternatively require no text reading stage at all and map input documents immediately to desired outputs (OCR-free) (Sassioui et al., 2023). The system by (Appalaraju et al., 2021) is only end-to-end during the pre-training phase. Also, (Klaiman and Lehne, 2021) propose a system that can be designed in an end-to-end manner, however the authors chose an external OCR engine in their work. These approaches first appeared in 2019, but only became more frequent in 2022 and 2023. Since this is a relatively new research direction, corresponding systems are relatively rare in the overall search period. The same is true for the 12 identified generative KIE methods (Klaiman and Lehne, 2021; Powalski et al., 2021; Wang et al., 2021b; Cao et al., 2022a, b; Kim et al., 2022; Cao et al., 2023; Davis et al., 2023; Deng et al., 2023a; Dhouib et al., 2023; He et al., 2023b; Tang et al., 2023), which first appeared in 2021, but now see an increased interest. The work by (He et al., 2023b) differs in the sense that the authors use an out-of-the-box Large Language Model (LLM) such as ChatGPT and prompt the model in specific ways including few-shot prompting to support KIE tasks. This in particular involves demonstrations for output formatting, but also making the models layout-aware in order to utilize positional relationships obtained by OCR. This category of KIE methods is able to output arbitrary texts using autoregressive decoding mechanisms. Key advantages are, that corresponding systems are not necessarily influenced by faulty OCR extractions as they can generate OCR-free representations of the target words (Cao et al., 2022b). They are also more flexible, as they can be conditioned of different DU tasks using distinct textual prompts (Kim et al., 2022; Davis et al., 2023). More traditional methods, which use a classification head for the downstream task such as KIE, would on the other hand need to be retrained with a different classification head in order to perform a different task (Davis et al., 2023).
@@
Only 10 papers explicitly use domain knowledge (Palm et al., 2017; Gal et al., 2019; Palm et al., 2019; Majumder et al., 2020; Belhadj et al., 2021; Garncarek et al., 2021; Sarkhel and Nandi, 2021; Tata et al., 2021; Arroyo et al., 2022; Du et al., 2022). (Oral et al., 2020) also consider it, however only regarding their RE part of the approach and not the subsystem that is used for KIE. The integration of domain knowledge mostly consists of hand-crafted input features, as discussed before. Only two approaches have been evaluated in real-world industry settings, namely (Oral et al., 2020) and (Hwang et al., 2021a). More precisely, the authors show the impact of the developed KIE methods on real-world document processing tasks. 14 of the analyzed papers represent an evolution of an existing approach. A prominent example is the LayoutLM family. Based on the original LayoutLM, many subsequent refinements and improvements in various aspects have been proposed (Xu et al., 2020, 2021; Gu et al., 2022; Huang et al., 2022; Shi et al., 2023). Another example is Chargrid (Katti et al., 2018), which has been the basis for most other grid-based approaches (see section 5.2.3). One could argue that 14 out of 96 papers being an evolution is a relatively high proportion, and that as a result a lot of research is being done on iterative refinements and, to a lesser extent, disruptive innovations that could achieve significantly better KIE results. 8 of the analyzed approaches are implemented using weakly-annotated data. In most cases, this means that besides the document images, only the textual target values are required (Palm et al., 2019; Garncarek et al., 2021; Klaiman and Lehne, 2021; Wang et al., 2021b). Some other approaches only require annotations on segment-level or word-level (Ning et al., 2021; Cao et al., 2022b). Corresponding approaches therefore do not need to be trained on fully annotated documents, which usually involves a labeling of each word including its bounding box and entity.

### v2

@@ Abstract.
Extracting key information from documents represents a large portion of business workloads and therefore offers a high potential for efficiency improvements and process automation. With recent advances in Deep Learning, a plethora of Deep Learning based approaches for Key Information Extraction have been proposed under the umbrella term Document Understanding that enable the processing of complex business documents. The goal of this systematic literature review is an in-depth analysis of existing approaches in this domain and the identification of opportunities for further research. To this end, 130 approaches published between 2017 and 2024 are analyzed in this study.
@@
copyright: rightsretainedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: CSURccs: Applied computing Document analysisccs: Applied computing Business process management systemsccs: Computing methodologies Information extractionccs: Computing methodologies Neural networks
@@ 1. Introduction
Complex business documents pose a significant challenge to KIE systems because they cannot be understood and processed as linear text sequences, as has been the case in most traditional KIE applications (Borchmann et al., 2021). These documents typically contain implicit and explicit cues and complex positional dependencies between certain text segments. In this regard, related research investigates different methods to integrate such cues into the model architectures in order to achieve better extraction results. Business documents also have special characteristics resulting from their connection to business processes. For example, different documents that are being processed as part of the same process run usually contain reoccurring information. These aspects and how they can improve corresponding KIE systems are worth investigating. In general, a process-oriented understanding is critical in order to adequately address the challenge of KIE in real-world contexts.
@@
    A SLR based on 130 approaches to provide a concise overview of DL-based KIE methods for business documents.
@@
The remainder of this manuscript is structured as follows: Chapter two provides background information on key concepts and nomenclature that are relevant to this SLR. In chapter three, we discuss related work in terms of existing surveys and illustrate, how this study differs. The methodology used for this SLR, and more specifically how relevant literature was identified, is illustrated in chapter four. Chapter five provides an overview of the results of the in-depth analysis. Chapter six serves as a discussion of the key findings and is dedicated to specific aspects of the identified literature. Based on the previous two chapters, we propose a research agenda and starting points for follow-up research in chapter seven. Chapter eight concludes the manuscript with a summary.
@@ 2.1. A business process perspective
In addition to implicit knowledge about predefined entities, the consideration of associated business processes can also be useful in the sense that the information flow between individual process steps could be taken into account. For example, one could consider how the data extracted from a business document is further processed in subsequent process steps. From this, conclusions could be drawn as to how corresponding business documents should be processed. One could also analyze whether – in addition to the documents themselves – there is an additional data flow between the document processing steps that could facilitate the understanding of the documents.
@@ 2.2. Document Understanding
Document Understanding (DU), based on concepts from Natural Language Processing (NLP) and Computer Vision (CV), is an umbrella term covering a wide range of document processing tasks, including KIE, table understanding, document layout analysis (DLA), document classification and visual question answering (VQA) (Borchmann et al., 2021). More recently, novel tasks have been proposed such as ”Key Information Localization and Extraction”, which is an extension of KIE that emphasizes the need to localize key information, e.g., by identifying bounding boxes [^1] A bounding box indicates the coordinates of an element in the document image by its top left corner as well as its width and height. (Skalický et al., 2022). For a brief introduction to DU tasks besides KIE that are not covered in this review, we refer to (Cui et al., 2021). There are also other terms that are used for this research area, such as Document Analysis, Document AI or Document Intelligence (Motahari et al., 2021; Cui et al., 2021).
@@ 2.4. Input documents
The term ”business document” appears less frequently in related literature and there is also no universal definition of such documents (Cristani et al., 2018). In general, business documents contain process-relevant details related to the internal and external operations of an organization, as these documents represent a central means of communication (Cui et al., 2021; Skalický et al., 2022). Business documents, like VRDs, pose a significant challenge to automated DU systems for a variety of reasons, as discussed by (Motahari et al., 2021). For example, corresponding documents exist in a variety of different formats and are often only available in scanned form due to their paper-based distribution. A scanned document is reduced to its image data and therefore requires techniques such as Optical Character Recognition (OCR) to make the content machine-readable (Oussaid et al., 2022). The layout and overall content of business documents can vary from highly structured to highly unstructured. Furthermore, business documents can have relationships to other documents and/or may consist of multiple documents in a hierarchical fashion. To understand business documents, it is therefore necessary to observe such interrelationships and to understand the temporal relationships in processes.
@@ 3. Existing reviews on Key Information Extraction
In the following, we discuss the most closely related work. (Subramani et al., 2020) provide a brief KIE section that is limited to discussions on a few relevant aspects. The survey also does not include an in-depth analysis of the approaches. In (Baviskar et al., 2021), the authors provide a comprehensive overview of document processing with many different facets. However, it offers limited discussion about KIE. Much refers specifically to NER, where the overall workflow as well as different methods regarding pre-processing and feature extraction are being discussed. The authors discuss some model architectures. Since their chosen search period includes work between 2010 and 2020, a large number of analyzed papers also do not necessarily use DL methods and/or consider VRDs. (Cui et al., 2021) cover a wide range of DL-based DU topics. KIE, on the other hand, is discussed relatively briefly. The survey discusses aspects related to transformer-based document processing methods from a general point of view. However, details about the employed model architectures and other KIE paradigms are not covered in detail. The survey of (Antonio et al., 2022) focuses on preliminary DU tasks such as text detection and text recognition. KIE is also covered, although only a small number of approaches are included. The authors present some individual approaches in detail, however there is a lack of comprehensive overview. In (Martínez-Rojas et al., 2023), document processing is positioned in the context of Robotic Process Automation. Although an extensive literature survey is conducted, there is no in-depth technical discussion of KIE approaches for VRDs. An overview of key methods for KIE is provided by (Li et al., 2023a), albeit at a relatively high level. The authors however provide an extensive outlook on future work. Besides KIE, (Sassioui et al., 2023) also cover VQA and document classification. The authors propose a taxonomy for DU along different dimensions and focus on KIE benchmark datasets. Some challenges as well as future work are also discussed. (Abdallah et al., 2024) present a review of transformer-based methods for DU tasks. They highlight key paradigms of corresponding models and showcase a few approaches in detail. A major focus of the review is the description of benchmark datasets and related performance comparisons. In general, the survey chooses to discuss a few transformer-based approaches in detail, but does not present a broad overview of DL-based approaches with respect to KIE.
@@
Overall, existing surveys either examine KIE from a very broad perspective, cover only a few approaches in detail, or alternatively include a higher number of published work, but at the expense of a detailed discussion of the underlying methods. This work on the other hand provides an in-depth analysis, both quantitatively and qualitatively. One key distinguishing factor is also the adoption of a business process perspective. As discussed in section 2.1, the consideration of business processes as well as general domain knowledge is of high importance for KIE systems. To this end, only the review by (Martínez-Rojas et al., 2023) also adopts a practice-oriented perspective during the analysis. Table 1 summarizes the differentiation of the previously mentioned related work along certain distinguishing factors.
@@ 4.1. Research questions
The first five research questions cover an in-depth analysis of the identified approaches, with the aim of examining the proposed methods and how they differ from each other. This includes key aspects such as input modalities, model architectures and data bases. Based on this, research questions six to eight adopt a more aggregated view and aim to identify the state of the art and derive recommendations for follow-up research.
@@ 4.2. Search procedure
The following six databases were used to obtain relevant literature: ACM Digital Library (ACM), ACL Anthology (ACL), AIS eLibrary (AIS), IEEE Xplore (IEEE), ScienceDirect (SD) and SpringerLink (SL). The search strings were carefully designed to include the commonly used terms for the research area as well as relevant keywords for the target domain (VRDs) and the application of DL architectures. If supported by the search engine, we also included terms that should not appear to filter out papers that are outside the scope of this work. The search period was limited to the range 2017 to 2024. Considering 2017 as the lower limit is a significant help to avoid false positive results, as KIE related literature typically did not use DL methods before 2017. The final search strings for each database, including the number of results at the time the query was run, can be found in table ST1 of the electronic supplementary material.
@@
  • EX3: The work does not apply the approach to VRDs, but to other domains and/or considers text-only inputs.
@@
The overall search process according to the PRISMA (Page et al., 2021) is visualized in figure SF2 of the electronic supplementary material. A total of 130 relevant approaches were identified for this SLR. A list of the identified approaches as well as their numbering, which is also used as a reference in this manuscript, can be found in table ST2 of the electronic supplementary material.
@@ 5.1. Overview
First, we provide a high-level overview of the analyzed work. To this end, table ST3 of the electronic supplementary material shows the results of analysis along different properties, each of which considers different aspects, also with practical applications in mind. Figure 3(a) visualizes the overall distribution of the KIE paradigms. Almost half of the approaches belong to the group of sequence-based methods. Two further common categories are graph-based systems and approaches that combine graph-based and sequence-based concepts. The grid representation is not as widely used as the graph representation, as only nine approaches are purely based on this paradigm. Large Language Model (LLM) based systems represent an emerging category with 11 identified approaches. Also, only seven of the analyzed papers cannot be assigned to any of the groups, which underlines the dominance of the previously mentioned paradigms for KIE systems. Interestingly, only the work by (Wang et al., 2021b) combines grid-based and sequence-based concepts into one KIE approach. Based on the distribution of the methods over time, as visualized in figure 3(b), one can see an increased interest in KIE research, especially since 2021. It is also noticeable that the dominance of sequence-based approaches first started in 2021, while graph-based methods were the most common group of approaches before that. The increased popularity of sequence-based methods in this period can probably be attributed to the influential work by (Xu et al., 2020), which was among the first to show state-of-the-art results using a transformer architecture and extensive pre-training procedures, which in turn led to a lot of follow-up research in this direction. There is also a large increase in graph-based methods in 2023 compared to previous years and compared to the other categories. This category has therefore remained relatively popular over time. Overall, 2024 had the highest number of papers published, showing that there is still a high level of interest in this area of research. In 2022, however, there was a considerable decrease in the number of corresponding KIE approaches. One reason for this observation may be the aforementioned increased interest in sequence-based methods, which typically require extensive pre-training. Therefore, many scholars spent 2022 on developing corresponding methods and on academic write-up, leading to a publication some time later in 2023 or 2024. Given the emergence of LLMs, they also start to appear in KIE research in 2023, with a significant increase in 2024, already becoming the second most common category.
@@
The analysis of the integrated input modalities shows that the vast majority of approaches integrate at least textual and layout-oriented features. This is not surprising, since these elements contain the most relevant information for document processing. Layout modalities are usually derived from bounding boxes, i.e., the coordinates of words in document images. Around 60% of the analyzed approaches integrate visual features obtained from the document images. As discussed in section 2.4, the integration of such visual cues into KIE systems can be crucial for a proper understanding of complex VRDs. It could be argued that – given this relevance – there are relatively few approaches that integrate image modalities. Nevertheless, the results show that visual cues have been increasingly integrated into the models over time, especially since 2021. In this regard, (Oral and Eryiğit, 2022) extensively investigate different methods to obtain and fuse visual modalities into KIE systems. They show that the choice of how visual information is represented depends on the underlying VRDs and that attention-based fusion mechanisms can outperform more basic methods such as concatenations of features vectors. In general, integrating multiple modalities can play an important role for DU systems, as they incorporate cues from different aspects that complement each other. For example, the image modality can provide relevant insights for documents with otherwise limited amounts of text and bounding boxes, while the layout modality is crucial for documents with very complex layouts and distinct structures (Yang et al., 2024a). Some approaches use hand-crafted input modalities. In most of these cases, they are integrated through custom input features such as boolean flags, whether a word in the document represents specific entities such as dates or monetary values. Another feature type is information about the position of a particular word with respect to the overall reading order. In some papers, hand-crafted inputs are based on syntactic features such as the numbers of characters of a word, details regarding the fonts or encoded character representations. The use of hand-crafted features is also often associated with grid-based approaches. Such methods usually define a custom encoding function that assigns a specific value to each element of the grid. For example, (Palm et al., 2017) assign a constant value to each pixel in a document image, such as an integer index of a character at each position – or the value 0 for blank parts of the document image. None of the approaches integrate meta data resulting from related business workflows as discussed in section 2.1, therefore lacking a practice-oriented perspective in this regard. Overall, hand-crafted features provide additional information beyond data that is explicitly contained in documents and can thus improve KIE systems. It must be said, however, that hand-crafted input features can potentially lead to a considerable amount of additional work in terms of labor-intensive data annotation (Saifullah et al., 2023a), which could also explain the less frequent use of this feature type. One could also argue that the inclusion of such hand-crafted features can potentially result in overfitting to the domain of the training documents – leading to poorer generalization capabilities across different domains. However, there is no work that analyzes this trade-off between integrating hand-crafted input features and the generalization capabilities across different domains.
@@
In terms of the underlying data basis, the median number of employed documents is 2,172. This is a relatively low amount, especially since the data basis usually has to be split into training and evaluation partitions. As a result, many of the approaches are often only trained on a small document corpus and thus probably with little variety in terms of layouts. Since DL-based DU models usually require large amounts of data for training (Saifullah et al., 2023a), it is debatable whether the proposed approaches have reached their full potential. Note that these observations do not include the data basis that is being used for pre-training purposes, which will be discussed in the context of sequence-based approaches in section 5.2.1. Deviating from the average, (Palm et al., 2019; Klaiman and Lehne, 2021; Sarkhel and Nandi, 2021) employ more than one million documents for implementing their KIE system. On the other hand, in some cases, only 199 documents are used for developing the models, while still achieving promising extraction results. This indicates that there are major differences in terms of data requirements of the individual approaches and that some architectures can manage with significantly less data. The vast majority of considered document types are receipts and forms, which are used by at least half of the approaches. The main reason for this is the fact that the most frequently used benchmark datasets are based on these document types. Besides, invoices are most often considered when authors do not use public benchmarks, but rather private in-house datasets.
@@
Regarding the aspect of reproducibility, implemented code is available for 50 approaches. In most cases, authors share their code directly, however there are exceptions where implementations have been made available by external sources through re-implementation. Model weights are even less commonly shared, although this would be particularly helpful for pre-trained models, as improvements and optimizations could be made by external parties on this basis. When either code or model weights are shared, they are mostly provided with a license that allows for commercial use. This is beneficial for organizations that want to integrate corresponding KIE systems into their own business processes, which in turn helps to disseminate the research efforts.
@@
A few approaches are independent of external OCR engines and are therefore either responsible both for text reading and information extraction (end-to-end) or alternatively require no text reading stage at all and map input VRDs immediately to desired outputs (OCR-free) (Sassioui et al., 2023). The system by (Appalaraju et al., 2021) is only end-to-end during the pre-training phase. Also, (Klaiman and Lehne, 2021) propose a system that can be designed in an end-to-end manner, however the authors chose an external OCR engine in their work. These approaches first appeared in 2019, but became more frequent since 2022. The same is true for the 23 identified generative KIE methods, which first appeared in 2021, but now see an increased interest, especially given the rise of generative LLMs. This category of KIE methods is able to output arbitrary texts using autoregressive decoding mechanisms. Key advantages are, that corresponding systems are not necessarily influenced by faulty OCR extractions as they can generate OCR-free representations of the target words (Cao et al., 2022b). They can also be more easily adapted to different DU tasks, for example by using distinct textual prompts (Kim et al., 2022; Davis et al., 2023). More traditional KIE methods that use a classification head would on the other hand need to be retrained with a different classification layer in order to perform a different task like document classification (Davis et al., 2023).
@@
Only 10 papers explicitly use domain knowledge. The integration of domain knowledge mostly consists of hand-crafted input features, as discussed before. This again emphasizes the lack of a practical perspective in related literature, as corresponding KIE systems do not consider insights that can be obtained from real-world workflows. Also, only three approaches have been evaluated in real-world industry settings. In these cases, the authors show the impact of the developed KIE methods on real-world document processing tasks – for example in terms of efficiency improvements. Therefore, there is a severe gap between research advancements and real-world applications as the majority of proposed models are not evaluated in practical settings. It is therefore possible that innovations that improve benchmarks may not be immediately suitable for real-world scenarios, e.g., due to extensive data requirements. 16 of the analyzed papers represent an evolution of an existing approach. A prominent example is the LayoutLM family. Based on the original LayoutLM (Xu et al., 2020), many subsequent refinements and improvements in various aspects have been proposed. Another example is Chargrid (Katti et al., 2018), which has been the basis for many other grid-based approaches (see section 5.2.3). A few of the analyzed approaches are implemented using weakly-annotated data. In most cases, this means that besides the document images, only the textual target values are required. Some other approaches only require annotations on segment-level or word-level. Corresponding approaches therefore do not need to be trained on fully annotated documents, which usually involves a labeling of each word including its bounding box and entity. Such approaches can be more suitable for real-world applications where data annotation is costly.